1)    Neon lights on the U of A campus are replaced at the rate of 200units per day. The physical plant orders the neon lights periodically. It costs $100 to initiate a purchase order. A neon light kept in storage is estimated to cost about $0.04 per day. The lead time between placing and receiving an order is 15 days. Determine the optimal inventory policy for ordering the neon lights.          



1
Expert's answer
2021-07-20T12:30:02-0400

D = Demand per day = 100 (the replacement per day)

S = Ordering cost = $100

h = Holding cost per item per day = $0.02

L = Average lead time = 12

We will apply EOQ formula to determine the optimal order quantity Q*.

"Q^*= \sqrt{\frac{2*D*S}{H}}=1000"

Reorder point = R = Lead time demand = "D*L = 100*12 = 1200" .

So, the optimal inventory policy is (Q,R) policy where Q = 1000 and R = 1200 i.e. an order of quantity equal to 1000 will be placed when the inventory in depleted up to a level of 1200.